# VPRTempo: A Fast Temporally Encoded Spiking Neural Network for Visual Place Recognition

### v1

@@ I Introduction
Spiking neural networks (SNNs) are computational tools that model how neurons in the brain send and receive information [64]. SNNs have attracted significant interest due to their energy efficiency, low-latency data processing, and deployability on neuromorphic hardware such as Intel’s Loihi or SynSense’s Speck [42, 41]. However, modeling the complexity of biological neurons limits the computational efficiency of SNNs, especially where resource-constrained systems with real-time applications are concerned. Using a SNN system that trades biological realism for overall system efficiency would be well suited for a variety of robotics tasks [1, 61, 57, 58].
@@
Figure 1: A-i Traversal image sequences from standard VPR datasets (Nordland, Oxford RobotCar) are filtered and processed to be converted into A-ii spikes where the pixel intensity determines amplitude. In order to temporally encode spikes to an abstracted theta oscillation, A-iii amplitudes determine the spike timing during a timestep. B-i Once spikes have been generated, they are passed into a SNN with 3 layers; an input, feature layer, and a one-hot encoded output layer where each output neuron represents one place. B-ii In order to scale the system for large datasets, we train individual expert module SNNs of up to 1000 places from subsets of an entire traversal dataset.
@@
One such task that could benefit from fast and efficient SNNs is the visual place recognition (VPR) problem, where incoming query images are matched to a potentially very large reference database, with a range of applications in robot localization and navigation [21, 22, 50, 5, 65, 59, 30, 3, 49], including as a loop closure component in Simultaneous Localization and Mapping (SLAM) [12, 23, 50, 63, 62]. However, VPR remains challenging as query images often have significant appearance changes when compared to the reference images, with factors such as time of day, seasonal changes, and weather contributing to these differences [66, 35, 33]. SNNs represent a unique way to solve VPR tasks for their low-latency and energy efficiency, especially where this is an important consideration for the overall design of a robot [26]. The flexible and adaptive learning that SNNs take advantage of make them ideal for VPR tasks, as network connections and weights will learn unique place features from traversal datasets.
@@
SNNs traditionally use a rate encoding method with substantial computational costs due to its reliance on spike counts [13, 17, 46]. Our work instead adopts a temporal encoding strategy inspired by central information storage techniques used by the brain, reducing the overall content the network processes during learning [52, 54]. This takes a different approach to what information each spike is representing and transmitting when compared to previous studies like those by Hussaini et al. [21, 22]. We leverage a simplified SNN framework [54], modular networks [6, 18, 25], and one-hot encoding for supervised training to substantially decrease training and inference times. Compared with previous systems, our SNN achieves query speeds exceeding 50 Hz on CPU hardware for large reference databases ($>$27k places) and even higher inference speeds when deployed on GPUs (as high as 500 Hz), indicating its potential usefulness for real-time applications [21, 36, 19, 5, 30, 32, 39, 65].
@@
    We present VPRTempo, a novel SNN system for VPR that for the first time encodes place information via a temporal spiking code (Figure 1), significantly increasing the information content of each spike.
@@
    We significantly lower the training time of spiking networks to sub-hour time frames and increase query speeds to real-time capability on both CPUs and GPUs, with the potential to represent tens of thousands of places even in resource-limited compute scenarios.
@@ II-A Visual place recognition
Several systems have been deployed to tackle this problem, ranging from simple approaches such as Sum of Absolute Differences (SAD) [36], over handcrafted systems like DenseVLAD [59] to learned approaches like NetVLAD [36, 5] and many others including [30, 3, 65]. These systems are commonly used in localization tasks such as loop closure in Simultaneous Localization And Mapping (SLAM) [60].
@@ II-C Spiking neural networks for robot localization
Neuromorphic computing involves the development of hardware, algorithms, and sensors inspired by neuroscience to solve a variety tasks [65, 55, 56, 61, 9]. Clearly, the brain is robustly capable of place recognition in a variety of contexts, providing a solid rationale to explore neuromorphic systems for VPR related challenges. A recent study exhibited prolonged training and querying times in an SNN developed with the Brian2 simulator, due to the reliance on rate encoding and modeling complex neuronal dynamics [21]. Robotic localization tasks have also employed SNNs to perform SLAM [14, 56], encode a contextual cue driven multi-modal hybrid sensory system for place recognition [65], and navigation [31, 16]. Our approach mitigates traditional barriers to deploying SNNs for real-time compute scenarios, enhancing the learning speed substantially by simplifying network processes and incorporating temporal spike encoding.
@@ III Methodology
This section is structured as follows: We first briefly introduce the underlying BliTNet in Section III-A. This is followed by the modular organization of networks in Section III-B. We then describe our novel efficient parallelizable training and inference procedure in Section III-C. After which we describe our implementation in Section III-D, the datasets used in Section III-E, evaluation metrics in Section III-F, the baseline methods that we compare and contrast our network to in Section III-G, and finally our strategy for optimizing system hyperparameters in Section III-H.
@@ III-A Temporal coding for visual place recognition
Network architecture: Each network consists of 3 layers. The input layer $L_{I}$ has as many neurons as the number of pixels in the input image, i.e. $W\cdot H$ neurons with $W$ and $H$ being the width and height of the input image, respectively. Neurons in $L_{I}$ are sparsely connected to a feature layer $L_{F}$. The output layer $L_{O}$ is one-hot encoded and fully connected to the $L_{F}$ layer such that each neuron $n_{i}\in L_{O}$ represents one geographically distinct place from a set of traversal images $\mathcal{D}=\{p_{1},p_{2},...,p_{N}\}$, where the number of places $N$ equals the number of output neurons $|L_{O}|$. For training that uses multiple traversals over different days/times, the same output neuron is used to encode that place so the network can learn changes to environments.
Network connections: Neuronal connections between the different layers can be excitatory, inhibitory, or both and are determined probabilistically. Connections from $L_{I}\rightarrow L_{F}$ are sparse and determined by an excitatory and inhibitory connection probability, $P_{exc}$ and $P_{inh}$ respectively. Connections from $L_{F}\rightarrow L_{O}$ are fully connected, such that every neuron in $L_{O}$ receives both excitatory and inhibitory input. Excitatory and inhibitory connections are individually defined since inhibitory weights undergo additional normalization (refer to the Homeostasis subsection in III-A).
Postsynaptic spike calculation: The network state $x^{n}_{j}$ for neuron $j$ in layer $m$ evolves in the following way:
@@
where $x^{m}_{i}$ is input neuron $i$ in layer $m$, $\theta\in[0,\theta_{max}]$ is the neuron firing threshold, $C$ is a constant input, and $W^{+nm}_{ji}$ and $W^{-nm}_{ji}$ are the excitatory and inhibitory weights, respectively. $x$ is clipped in the range $[0,1]$ in case of negative or large spikes.
@@
where $W^{nm}_{ij}$ refers to both positive and negative weights, $\eta_{\text{STDP}}$ is the STDP learning rate, $f_{j}^{n}$ is the target firing rate of neuron $j$ in layer $n$, $\Theta(\cdot)$ is the Heaviside step function, $x$ is the spike amplitude of pre- or post-synaptic neurons, and $t$ is the timestep. Effectively, a pre-synaptic neuron firing before a post-synaptic one will undergo a positive weight shift (synaptic potentiation). The inverse of this scenario will result in a negative weight shift (synaptic depression). To ensure that the weights persist as positive or negative connections, they cannot change sign during training. Therefore, if a sign change is detected then that connection will be reset to $\epsilon=10^{\pm 6}$.
@@
$\begin{gathered}W^{-nm}_{ji}(t)\leftarrow W^{-nm}_{ji}\Big{(}1-\eta_{\text{STDP}}(t)\cdot\Theta\big{(}\sum_{i}x^{m}_{i}(t)\big{)}\Big{)}.\end{gathered}$ (9)
When the network is first initialized, neurons have randomly generated firing thresholds which may never be crossed depending on the input. To prevent neurons from being consistently inactive, we use an Intrinsic Threshold Plasticity (ITP) learning rate $\eta_{\text{ITP}}$ to adjust the spiking threshold to an ideal value for post-synaptic neurons:
@@ III-B Modular place representation for temporal codes
The brain encodes scene memory with place cells into sparse circuitry using engrams, individual circuits of neurons that activate in response to external cues for recall [37]. To mimic the concept of an engram, we used modules of individually trained networks to learn subsets of traversal datasets since the brain does not encode place recognition in a singular network.
By training multiple individual networks we gain several advantages: 1) Smaller networks train faster and are more accurate, 2) since the probability of connections, firing thresholds, and firing rates across $L_{I}\rightarrow L_{F}$ is uniformly random, a query image $q$ trained in network $N_{i}$ is unlikely to activate output neurons with a high enough amplitude in other networks to generate a false positive match, and 3) similar to what has been shown in previous work [21], modularizing greatly improves the scalability of the system to learn more places for VPR.
@@
$\begin{gathered}U=\underset{i=\{1,...,|N|\}}{\cup}\ N_{i}\ \text{with}\ N_{i}\ \cap N_{j}\ =\ \emptyset\ \forall_{i}\ \neq\ j.\end{gathered}$ (16)
@@ III-C Efficient implementation
Our efficient implementation for VPRTempo involves training and querying a 3D weight tensor $\mathbf{T}\in\mathbb{R}^{|N|*|L_{i}|*|L_{j}|}$ with $|N|$ being the number of modules, $|L_{i}|$ the number of the neurons in the pre-synaptic layer (either $L_{I}$ or $L_{F}$), and $L_{j}$ being the number of neurons in the post-synaptic layer ($L_{F}$ or $L_{O}$). This setup capitalizes on parallel computing to boost efficiency and speed [43]. During training, the weight tensor $\mathbf{T}$ is multiplied by an image tensor $\mathbf{I}\in\mathbb{R}^{|N|*1*|L_{i}|}$ that holds images associated with a particular module and their input spike rates. At deployment time, the single query image is fed to all modules concurrently, leveraging parallel processing to decrease the overall inference time.
III-D Implementation details
Our network was developed and implemented in Python3 and PyTorch [43]. The system is trained on non-overlapping and geographically distinct places from multiple standard traversal datasets (Nordland and Oxford RobotCar [40, 34]), as commonly used in the literature [21, 22, 50, 38]. All reference and query input images underwent pre-processing for gamma correction, resizing, and patch normalization [36]. Specifically, Gamma correction was used to normalize pixels $\rho_{i}$ of the input images:
@@ III-D Implementation details
where $\gamma=\frac{e^{\lambda\times 255}}{e^{\mu}}$ with $\lambda=0.5$ and $\mu=\bar{\rho_{i}}$. Normalized images were then resized to $W\times H=28\times 28$ pixels and patch-normalized with patch sizes $W_{P}\times H_{P}=7\times$ 7 pixels [36]. Network spikes $x$ are defined as floating point amplitudes in the range [0, 1], where $x=1$ is a full spike and $x=0$ is no spike. Initial input spikes are generated from pixel intensity values $i=[0,255]$ from training or test images and converted to amplitudes by $x=\frac{i}{255}$.
As an abstraction of theta oscillations in the brain, the spike amplitude determines where in a timestep the spike occurs, where 1 timestep is equivalent to one phase of theta (Figure 1). The network hyperparameters after performing a grid search (Section III-H) were set as to the values in Table I. Network modules were trained to learn 1100 places with 3 modules (totalling 3300 places) and 900 places with 3 modules (totalling 2700 places) for the comparisons presented in Figure 2 and Table II, Figure 3, respectively.
Each training epoch contained images from multiple traversals of the datasets of different day/time and were trained for 4 epochs in total. The training of each module is done simultaneously such that for each epoch, multiple unique places can be learned at the same time. After the network trained for sufficient epochs, all modules were simultaneously queried with independent network activation states to perform place matching.
III-E Datasets
@@
Images from both datasets were sub-sampled every 8 seconds, which is roughly 100 metres apart for Nordland, and 20 metres for Oxford RobotCar resulting in 3300 and 450 total places, respectively. It is important to note that while VPRSNN trained on these number of images, it only queried 2700 and 360 places respectively due to the 20% of input images used for calibration [21]; note also that our proposed VPRTempo does not require a calibration dataset.
III-F Evaluation metrics
When a query image $q$ is processed by the network, the matched place $\hat{p}$ is the place $i$ that is assigned to the neuron $n_{i}$ in the $L_{O}$ with the highest spike amplitude $f_{i}$:
$\begin{gathered}\hat{p}=\operatorname*{arg\,max}_{i}f_{i}.\end{gathered}$ (21)
Our main metric for overall evaluation of the network is precision at 100% recall (P@100R) [50], which was used to determine performance when compared to other state-of-the-art VPR systems. P@100R measures the percent of correct matches when the network is forced to match every query image with a reference image. Recall at $N$ (R@$N$) measures if the true place is within the top $N$ matches. A match is only considered a true positive if it aligns with the exact ground truth, i.e. there is no ground truth tolerance around the true match.

### v2

@@ I Introduction
Spiking neural networks (SNNs) are computational tools that model how neurons in the brain send and receive information [1]. SNNs have attracted significant interest due to their energy efficiency, low-latency data processing, especially when deployed on neuromorphic hardware such as Intel’s Loihi or SynSense’s Speck [2, 3]. However, modeling the complexity of biological neurons limits the computational efficiency of SNNs, especially where resource-constrained systems with real-time applications are concerned. Using a SNN system that trades biological realism for overall system efficiency would be well suited for a variety of robotics tasks [4, 5, 6, 7].
@@
Figure 1: A-i Image sequences from standard VPR datasets (Nordland, Oxford RobotCar) are filtered and processed to be converted into A-ii spikes where the pixel intensity determines amplitude. A-iii In order to temporally encode spikes to an abstracted theta oscillation, amplitudes determine the spike timing during a timestep. B-i Once spikes have been generated, they are passed into a SNN with 3 layers; an input, feature layer, and a one-hot encoded output layer where each output neuron represents one place. B-ii In order to scale the system for large datasets, we train individual expert module SNNs of up to 1000 places from subsets of an entire reference dataset.
@@
One such task that could benefit from fast and efficient SNNs is visual place recognition (VPR), where incoming query images are matched to a potentially very large reference database, with a range of applications in robot localization and navigation [8, 9, 10, 11, 12, 13, 14, 15, 16], including as a loop closure component in Simultaneous Localization and Mapping (SLAM) [17, 18, 10, 19, 20]. However, VPR remains challenging as query images often have significant appearance changes when compared to the reference images, with factors such as time of day, seasonal changes, and weather contributing to these differences [21, 22, 23]. SNNs represent a unique way to perform VPR for their low-latency and energy efficiency, especially where this is an important consideration for the overall design of a robot [24]. The flexible and adaptive learning that SNNs take advantage of make them ideal for VPR, as network connections and weights will learn unique place features from reference datasets.
@@
SNNs employ a variety of encoding mechanisms, with the traditional rate encoding suffering from substantial computational costs due to its reliance on spike counts [25, 26, 27]. Our work instead adopts a temporal encoding strategy inspired by central information storage techniques used by the brain, reducing the overall content the network processes during learning [28, 29]. This takes a different approach to what information each spike is representing and transmitting when compared to previous studies like those by Hussaini et al. [8, 9]. We leverage a simplified SNN framework [29], modular networks [30, 31, 32], and one-hot encoding for supervised training to substantially decrease training and inference times. Compared with previous systems, our SNN achieves query speeds exceeding 50 Hz on CPU hardware for large reference databases ($>$27k places) and even higher inference speeds when deployed on GPUs (as high as 500 Hz), indicating its potential usefulness for real-time applications [8, 33, 34, 11, 14, 35, 36, 12].
@@
    We present VPRTempo, a novel SNN system for VPR that, to the best of our knowledge, for the first time encodes place information via a temporal spiking code (Figure 1), significantly increasing the information content of each spike.
@@
    We significantly lower the training time of spiking networks to under an hour and increase query speeds to real-time capability on both CPUs and GPUs, with the potential to represent tens of thousands of places even in resource-limited compute scenarios.
@@ II-A Visual place recognition
Several systems have been deployed to tackle this problem, ranging from simple approaches such as Sum of Absolute Differences (SAD) [33], over handcrafted systems like DenseVLAD [13] to learned approaches like NetVLAD [33, 11] and many others including [14, 15, 12]. These systems are commonly used in localization tasks such as loop closure in Simultaneous Localization And Mapping (SLAM) [41]. Although these approaches have shown to be highly effective at accurate place recognition, they require computational and time intensive training regimes and learn generalized feature extraction, rather than actually learning the specific location.
@@ II-C Spiking neural networks for robot localization
Neuromorphic computing involves the development of hardware, algorithms, and sensors inspired by neuroscience to solve a variety tasks [12, 47, 48, 5, 49]. Clearly, the brain is robustly capable of place recognition in a variety of contexts, providing a solid rationale to explore neuromorphic systems for VPR related challenges. A recent study exhibited prolonged training and querying times in an SNN, due to the reliance on the low spike efficiency of rate encoding and modeling complex neuronal dynamics [8]. Robotic localization tasks have also employed SNNs to perform SLAM [50, 48], encode a contextual cue driven multi-modal hybrid sensory system for place recognition [12], and navigation [51, 52]. Our approach mitigates traditional barriers to deploying SNNs for real-time compute scenarios, enhancing the learning speed substantially by simplifying network processes and incorporating temporal spike encoding.
@@ III Methodology
This section is structured as follows: We first briefly introduce the underlying BliTNet in Section III-A. This is followed by the modular organization of networks in Section III-B. We then describe our novel efficient parallelizable training and inference procedure in Section III-C.
@@ III-A Temporal coding for visual place recognition
Network architecture: Each network consists of 3 layers. The input layer $L_{I}$ has as many neurons as the number of pixels in the input image, i.e. $W\cdot H$ neurons with $W$ and $H$ being the width and height of the input image, respectively. Neurons in $L_{I}$ are sparsely connected to a feature layer $L_{F}$. The $L_{F}$ layer is fully connected to the output layer $L_{O}$, which is one-hot encoded, such that each neuron $n_{i}\in L_{O}$ corresponds to one geographically distinct place from a set of reference images $\mathcal{D}=\{p_{1},p_{2},...,p_{N}\}$. The number of output neurons $|L_{O}|$ matches the number of distinct places $N$. For training that uses multiple traversals over different days/times, the same output neuron is used to encode that place so the network can learn changes in environments.
Network connections: Connections from $L_{I}\rightarrow L_{F}$ are sparse and determined by excitatory and inhibitory connection probabilities $P_{exc}$ and $P_{inh}$ respectively. Connections from $L_{F}\rightarrow L_{O}$ are fully connected, such that every neuron in $L_{O}$ receives both excitatory and inhibitory input. Excitatory and inhibitory connections are individually defined since inhibitory weights undergo additional normalization (refer to the Homeostasis subsection below).
Neuron dynamics: The neuron state $x^{n}_{j}$ for neuron $j$ in layer $n$ evolves in the following way:
@@
where $x^{m}_{i}$ is input neuron $i$ in layer $m$, $\theta\in[0,\theta_{max}]$ is the neuron firing threshold, $C$ is a constant input, and $W^{+nm}_{ji}$ and $W^{-nm}_{ji}$ are the excitatory and inhibitory weights, respectively. $x^{n}_{j}$ is clipped in the range $[0,1]$ to prevent spike vanishing or exploding, respectively.
@@
where $W^{nm}_{ij}$ refers to both positive and negative weights, $\eta_{\text{STDP}}$ is the STDP learning rate, $f_{j}^{n}$ is the target firing rate of neuron $j$ in layer $n$, $\Theta(\cdot)$ is the Heaviside step function, $x^{m}_{i}$ and $x^{n}_{j}$ are the neuron states of pre- or post-synaptic neurons respectively, and $t$ is the timestep. Effectively, a pre-synaptic neuron firing before a post-synaptic one will undergo a positive weight shift (synaptic potentiation). The inverse of this scenario will result in a negative weight shift (synaptic depression). To ensure that the weights persist as positive or negative connections, they cannot change sign during training. Therefore, if a sign change is detected then that connection will be reset to be $\epsilon=10^{\pm 6}$, whereby the sign of $\epsilon$ matches the sign prior to the sign change.
@@
$\begin{gathered}\hat{W}^{-nm}_{ji}(t)\leftarrow W^{-nm}_{ji}(t)\Big{(}1-\eta_{\text{STDP}}(t)\cdot\Theta\big{(}\sum_{i}x^{m}_{i}(t)\big{)}\Big{)}.\end{gathered}$ (9)
When the network is first initialized, neurons have randomly generated firing thresholds which may never be crossed depending on the input. To prevent neurons from being consistently inactive, an Intrinsic Threshold Plasticity (ITP) learning rate $\eta_{\text{ITP}}$ is used to adjust the spiking threshold to an ideal value for post-synaptic neurons:
@@ III-B Modular place representation for temporal codes
The brain encodes scene memory with place cells into sparse circuitry using engrams, individual circuits of neurons that activate in response to external cues for recall [62]. To mimic the concept of an engram we train multiple expert networks, which provides several advantages: 1) Smaller networks train faster and are more accurate, 2) the heterogeneity of initial network seeds is uniformly random, thus a query image $q$ trained in network $N_{i}$ is unlikely to activate output neurons with a high enough amplitude in other networks to generate a false positive match, and 3) similar to previous work [8], modularizing improves system scalability to learn more places.
Choosing the number of expert modules to employ depends on the dataset and number of places being encoded. Accuracy is influenced by $T$ and learning rate annealment therefore too few or too many places will affect the systems learning capability.
@@
$\begin{gathered}U=\bigcup_{i=1}^{|N|}N_{i}\quad\text{with}\quad N_{i}\cap N_{j}=\emptyset\quad\forall i\neq j.\end{gathered}$ (16)
@@ III-C Efficient implementation
Our efficient implementation for VPRTempo involves training and querying a 3D weight tensor $\mathbf{T}\in\mathbb{R}^{|N|\cdot|L_{i}|\cdot|L_{j}|}$ with $|N|$ being the number of modules, $|L_{i}|$ the number of the neurons in the pre-synaptic layer (either $L_{I}$ or $L_{F}$), and $L_{j}$ being the number of neurons in the post-synaptic layer ($L_{F}$ or $L_{O}$). This setup capitalizes on parallel computing to boost efficiency and speed [60]. During training, the weight tensor $\mathbf{T}$ is multiplied by an image tensor $\mathbf{I}\in\mathbb{R}^{|N|\cdot 1\cdot|L_{i}|}$ that holds images associated with a particular module and their input spike rates. Single query images are fed to all modules concurrently, leveraging parallel processing to decrease the overall inference time.
IV Experimental Setup and Implementation
Here we describe our implementation in Section IV-A, the datasets used in Section IV-B, evaluation metrics in Section IV-C, the baseline methods that we compare and contrast our network to in Section IV-D, and finally our strategy for optimizing system hyperparameters in Section IV-E.
IV-A Data processing
Our network was developed and implemented in Python3 and PyTorch [60], converting the original BLiTNet implementation from NumPy to efficiently use multi-dimensional tensors and GPU acceleration for improved network performance. The system is trained on non-overlapping and geographically distinct places from multiple standard VPR datasets (Nordland and Oxford RobotCar [37, 38]), as commonly used in the literature [8, 9, 10, 63]. All reference and query input images underwent pre-processing for gamma correction, resizing, and patch normalization [33]. Specifically, Gamma correction was used to normalize pixels $\rho_{i}$ of the input images:
@@ IV-A Data processing
where $\gamma=\frac{e^{\lambda\times 255}}{e^{\mu}}$ with $\lambda=0.5$ and $\mu=\bar{\rho_{i}}$. Normalized images were then resized to $W\times H=28\times 28$ pixels and patch-normalized with patch sizes $W_{P}\times H_{P}=7\times$ 7 pixels [33]. Neuron states $x$ are defined as floating point amplitudes in the range [0, 1], where $x=1$ is a full spike and $x=0$ is no spike. Initial input spikes are generated from pixel intensity values $i=[0,255]$ from training or test images and converted to amplitudes by $x=\frac{i}{255}$.
As an abstraction of theta oscillations in the brain, the neuron state determines the timing of a spike within a timestep (Figure 1). The network hyperparameters after performing a grid search (Section IV-E) were set as to the values in Table I. Network modules were trained to learn 1100 places with 3 modules (totalling 3300 places) from the Nordland dataset for the comparisons presented in Figure 2 and Table II, Figure 3.
Training each module is done simultaneously such that for each epoch, multiple unique places can be learned at the same time. After the network trained for sufficient epochs, all modules were simultaneously queried with independent network activation states to perform place matching. The network was trained for a total of 4 epochs.
IV-B Datasets
@@
Images from both datasets were sub-sampled every 8 seconds, which is roughly 100 metres apart for Nordland, and 20 metres for Oxford RobotCar resulting in 3300 and 450 total places, respectively. It is important to note that while VPRSNN trained on these number of images, it reported accuracy performance on 2700 and 360 places respectively for the accuracy experiments due to the 20% of input images used for calibration [8]; note also that our proposed VPRTempo does not require a calibration dataset. For fairness of comparison, we present accuracy measurements that have omitted the first 20% of reference images.
IV-C Evaluation metrics
When a query image $q$ is processed by the network, the matched place $\hat{p}$ is the place $i$ that is assigned to the neuron $x_{i}$ in the $L_{O}$ with the highest neuron state $x_{i}$:
$\begin{gathered}\hat{p}=\operatorname*{arg\,max}_{i}x_{i}.\end{gathered}$ (21)
Network performance was measured with precision and recall [10]. Precision is the accuracy of identified places among all the predictions made, whereas recall is the coverage of true places successfully identified from all possible true places. Recall at $N=1$ ($R@1$) measures the percent of correct matches when the network is forced to match every query image with a reference image. Recall at $N$ ($R@N$) measures if the true place is within the top $N$ matches. A match is only considered a true positive if it aligns with the exact ground truth, i.e. no ground truth tolerance was employed.